Solve by completing the square
x^2+8x+13=0

Answer by stanbon(75887)

Solve by completing the square
x^2+8x+13=0
x^2+8x+? = -13+?
x^2+8x+4^2 = -13+4^2
(x+4)^2 = 3
x+4 = sqrt3 or x+4 = -sqrt3
x = -4+sqrt3 or x = -4-sqrt3
===============
Cheers,
Stan H.